Find an equation of the line parallel to the line with equation 4x - 2y = 8 and which passes through the point (2, -3)
Answer by lwsshak3(11628)

Find an equation of the line parallel to the line with equation 4x - 2y = 8 and which passes through the point (2, -3)
**
Equation for a straight line: y=mx+b, m=slope, b=y-intercept
4x-2y=8
2y=4x-8
y=2x-4
slope=2
..
slope of line parallel to given line=2
y=2x+b
solve for b using coordinates of given point(2,-3)
-3=2*2+b
b=-7
equation: y=2x-7